# Large Language Model Can Transcribe Speech in Multi-Talker Scenarios with Versatile Instructions thanks: *Equal Contribution

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have revolutionized various domains, bringing significant progress and new opportunities. Despite progress in speech-related tasks, LLMs have not been sufficiently explored in multi-talker scenarios. In this work, we present a pioneering effort to investigate the capability of LLMs in transcribing speech in multi-talker environments, following versatile instructions related to multi-talker automatic speech recognition (ASR), target talker ASR, and ASR based on specific talker attributes such as sex, occurrence order, language, and keyword spoken. Our approach utilizes WavLM and Whisper encoder to extract multi-faceted speech representations that are sensitive to speaker characteristics and semantic context. These representations are then fed into an LLM fine-tuned using LoRA, enabling the capabilities for speech comprehension and transcription. Comprehensive experiments reveal the promising performance of our proposed system, MT-LLM, in cocktail party scenarios, highlighting the potential of LLM to handle speech-related tasks based on user instructions in such complex settings [^1] The code, model, and samples is being organized and will be open-sourced in the camera-ready version..
@@ I Introduction
Large language models (LLMs) have experienced rapid and significant advancements recently, achieving or even surpassing human-level proficiency in numerous natural language processing (NLP) tasks [1, 2, 3]. These advancements have sparked interest in exploring the capabilities of LLM in multi-modal perception, including speech [4, 5, 6, 7], vision [1, 8, 9], and content generation [10, 11]. Several studies have investigated speech-related LLM, which typically involve a fine-tuned text LLM following speech-related instructions and pairing with auxiliary audio encoders [4, 5, 6, 7]. The audio encoder extracts acoustic representations and adapts them into the input feature space of LLM, enabling the LLM to perform various speech tasks such as automatic speech recognition (ASR), speech translation (ST), speaker verification (SV), and speech question answering (SQA), among others. However, despite the progress, the potential of speech LLMs in cocktail party scenarios—where multiple talkers speak simultaneously and overlapping occurs—has not yet been sufficiently exploited.
In recent year, various end-to-end approaches have garnered interest and been developed to tackle multi-talker ASR task, which involves simultaneously transcribing speech from multiple talkers. These studies are based on Permutation Invariant Training (PIT) [12, 13, 14], Heuristic Error Assignment Training (HEAT) [15, 16], or Serialized Output Training (SOT) [17, 18, 19, 20] to match predictions with corresponding target labels for loss calculation. However, these approaches typically transcribe speech from all talkers indiscriminately and fail to associate transcriptions with specific talkers, unless an additional external [21, 22] or internal [23, 24, 25] model is employed to extract speaker information. Although several studies [26, 27] proposed handling multi-talker ASR in conjunction with other tasks within a single model, the addressed tasks remain constrained and lack the flexibility to address various user requirements specifying talker attributes such as please transcribe the talker who said the word “strawberry”.
@@ II Methods
We propose empowering a text-based LLM to act as a versatile instruction follower for speech recognition in multi-talker speech scenarios. The proposed method consists of three major components: a large language model as the foundational model fine-tuned with LoRA, dual speech encoders with corresponding adapters, and training data construction. We denote the proposed model as MT-LLM in the subsequent sections.
@@ II-C Task Descriptions
To validate the performance of MT-LLM in executing speech recognition based on instructions in multi-talker scenarios, we simulated multi-talker overlapped speech audios from single-talker speech corpora, and designed six tasks along with corresponding instructions and generated the respective target text label. As illustrated in Fig. 1 (b), the text instructions [^3] The detailed instructions used for each task will be listed in the upcoming open-sourced repository. pertain to the following tasks:
@@ III-A Training Data Construction
We simulated multi-talker speech audios from single-talker speech corpora to accommodate the versatile task instructions, following the protocol outlined in [17]. Specifically, the utterances are simulated primarily from the 960-hour LibriSpeech [33] training set, comprising mixtures of two or three talkers. The start time for each talker is randomly sampled, resulting in overlapped mixtures. Additionally, the 180-hour German subset of CoVoST 2444Note that CoVoST 2 is originally designed for speech translation, while we use its German subset for ASR in our study. [34], along with its corresponding German target text, is employed to mix with LibriSpeech utterances to support the Target-Lingual ASR task.
@@ III-C Model Settings and Training details
The proposed MT-LLM employs the encoder of Whisper-large-v2555https://huggingface.co/openai/whisper-large-v2 and WavLM-base [^6] https://huggingface.co/microsoft/wavlm-base to extract speech representations, with Llama-2-chat-7b as the backbone LLM. All parameters of above models are frozen. The LLM is fine-tuned using LoRA with a rank of 32. The adapters for both speech encoders consist of two convolution layers to down-sample and align the representations within the temporal domain, followed by a bottleneck adapter [37] and a linear layer. The outputs of both adapters have a time stride of 80 ms and a dimension of 2,048. The total number of parameters in MT-LLM is 7.55 billion, of which 1% (76.6 million) are trainable.
@@ IV-A Results on Multi-Talker ASR
We evaluate the single-talker ASR performance on the LibriSpeech and CoVoST 2 German test set, and multi-talker ASR performance on the 2- and 3-speaker LibriSpeechMix and the home-made En-De-Mixed test set. As shown in Table I, the proposed MT-LLM demonstrates promising results across all datasets, outperforming D2V-Sidecar-DB [26], a recent representative work employing PIT. SOT-Conformer [23] is a state-of-the-art model specifically designed and extensively trained for multi-talker ASR on the LibriSpeechMix dataset. On the 2-speaker LibriSpeechMix, MT-LLM achieves results comparable to SOT-Conformer, while lags behind on the 3-speaker set. Given that MT-LLM serves as an exploratory work aimed at exploring the potential of LLM for versatile instruction-based ASR in complex environments, rather than solely pushing the limits of multi-talker ASR, we argue that the gap is understandable. Despite being trained on both single-talker and multi-talker data, we note that SOT-Conformer falls short in single-talker scenarios, indicating its specialization in multi-talker ASR impairs its single-talker performance. In contrast, MT-LLM perform consistently good in single-talker ASR, demonstrating that its instruction-understanding capabilities help maintain high performance in simpler, single-talker settings. SALMONN [5] is a speech LLM trained on various speech-related tasks including overlapped speech recognition. However, using its official prompt, we observed significant inferior performance compared to our approach as shown in Table I.
We list wav2vec2-Large-XLSR-53-German model [^7] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-german performance on De for reference. Given that the audio samples containing German is relatively limited and noiser than LibriSpeech, MT-LLM still demonstrates satisfactory results on De and En-De-Mix test sets, showcasing the model’s ability to simultaneously distinguish and transcribe multiple languages from multi-talker speech.
@@ IV-B Results on Versatile Tasks Based on Instructions
We investigate the performance of MT-LLM on various ASR tasks described in Section II-C. The evaluations are conducted using the 2-speaker and 3-speaker LibriSpeechMix and En-De-Mix test sets, with the results presented in Table II and Table III.
@@
The results in Table II and Table III highlight the effectiveness of the MT-LLM model across different instruction-based ASR tasks. In both the 2-speaker and 3-speaker mixed scenarios. MT-LLM achieves impressive performance for Target-Talker (TT), Keyword Tracing (KT), and Sex-Specific (SS) ASR tasks, which are within a reasonable range of the best-matching results. This suggests that MT-LLM is proficient at isolating and accurately transcribing speech from specified talkers based on instructions focusing on different speaker attributes. However, for Order-Specific (OS) ASR tasks, there is a gap compared to the best-matching result, indicating that MT-LLM slightly falls short in determining the order in which the talker appears. We anticipate significant improvements by using additional positional embedding for the speech part, employing larger speech modality adapters, or fully fine-tuning the speech encoder with heavier training as in [23]. Target-Lingual (TL) ASR task present a more pronounced challenge for the model, due to that the mixed audios containing German are relatively limited and noisier compared to English set. We observe that performance of the TL task in both English and German is close, resulting in similar WER to best-matching situation. This indicates room for improvement through language-specific embeddings or higher-quality multi-lingual data.
@@ IV-C Ablation Study
Fist, we conduct an ablation study to examine the impact of using dual speech encoders. We train two models specifically on the multi-talker (MT) ASR task: one using dual speech encoders and the other using only the Whisper encoder. As shown in Table IV, incorporating the WavLM encoder into the architecture leads to notable enhancements in performance, underscoring the importance of multi-layer acoustic information captured by WavLM in improving speech recognition accuracy, particularly in three-talker scenarios.
We also investigate the effect of multi-task training on improving the basic MT ASR task. As illustrated in Table IV, training on a variety of tasks not only endows the system with the versatility to handle diverse ASR tasks but also enhances performance on the basic MT ASR task. This indicates that the different tasks are interdependent and complementary, helping to supervise the model and enhance its overall speech comprehension capabilities.
@@ IV-D Limitations and Future Work
Despite MT-LLM’s promising performance on various ASR tasks in multi-talker scenarios, we acknowledge several limitations. First, the primary intention of this study is to exploring the ability of LLMs to capture specific talkers according to instructions, and transcribe their speech in multi-talker scenarios. Therefore, MT-LLM is not designed to be a universal speech LLM for a broader spectrum of tasks. Second, the experiments are conducted on limited simulated datasets for insightful observation, rather than on real-world datasets, due to resource constraints. In the future, we anticipate the development of a more comprehensive speech LLM for cocktail party environments with meticulously crafted training data and scheme.
@@ V Conclusion
In this work, we present a pioneering exploration into the use of large language models (LLMs) for instruction-based speech recognition in multi-talker scenarios. We utilize the Whisper encoder to extract semantic context information and WavLM to capture multi-layer acoustic information indicating speaker characteristics, thereby enabling the foundational LLM to effectively handle speech modality input. With parameter-efficient fine-tuning, our proposed system, MT-LLM, demonstrates remarkable capabilities in comprehending and transcribing speech based on versatile instructions related to multi-talker ASR, target talker ASR, and ASR based on specific talker attributes such as sex, occurrence order, language, and keyword spoken. Comprehensive experiments reveal promising performance in complex multi-talker environments, highlighting the potential of LLMs to enhance speech-related tasks and improve human-computer interaction in challenging settings.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have revolutionized various domains, bringing significant progress and new opportunities. Despite progress in speech-related tasks, LLMs have not been sufficiently explored in multi-talker scenarios. In this work, we present a pioneering effort to investigate the capability of LLMs in transcribing speech in multi-talker environments, following versatile instructions related to multi-talker automatic speech recognition (ASR), target talker ASR, and ASR based on specific talker attributes such as sex, occurrence order, language, and keyword spoken. Our approach utilizes WavLM and Whisper encoder to extract multi-faceted speech representations that are sensitive to speaker characteristics and semantic context. These representations are then fed into an LLM fine-tuned using LoRA, enabling the capabilities for speech comprehension and transcription. Comprehensive experiments reveal the promising performance of our proposed system, MT-LLM, in cocktail party scenarios, highlighting the potential of LLM to handle speech-related tasks based on user instructions in such complex settings [^1] The code, model, and samples are available at https://github.com/cuhealthybrains/MT-LLM.
@@ I Introduction
Large language models (LLMs) have experienced rapid and significant advancements recently, achieving or even surpassing human-level proficiency in numerous natural language processing (NLP) tasks [1, 2, 3]. These advancements have sparked interest in exploring the capabilities of LLM in multi-modal perception[4], including speech [5, 6, 7], vision [1, 8, 9], and content generation [10, 11, 12]. Several studies have investigated speech-related LLMs, which typically involve a fine-tuned text LLM following speech-related instructions and pairing with auxiliary audio encoders [5, 6, 7]. The audio encoder extracts acoustic representations and adapts them to the input feature space of LLM, allowing LLM to perform various speech tasks such as automatic speech recognition (ASR), speech translation (ST), speaker verification (SV), and speech question answering (SQA), among others. However, despite the progress, the potential of speech LLMs in cocktail party scenarios—where multiple talkers speak simultaneously and overlapping occurs—has not yet been sufficiently exploited.
In recent year, various end-to-end approaches have garnered interest and been developed to tackle multi-talker ASR task, which involves simultaneously transcribing speech from multiple talkers. These studies are based on Permutation Invariant Training (PIT) [13, 14, 15], Heuristic Error Assignment Training (HEAT) [16, 17], or Serialized Output Training (SOT) [18, 19, 20, 21, 22] to match predictions with corresponding target labels for loss calculation. However, these approaches typically transcribe speech from all talkers indiscriminately and fail to associate transcriptions with specific talkers, unless an additional external [23, 24] or internal [25, 26, 27] model is employed to extract speaker information. Although several studies [28, 29] proposed handling multi-talker ASR in conjunction with other tasks within a single model, the addressed tasks remain constrained and lack the flexibility to address various user requirements specifying talker attributes such as please transcribe the talker who said the word “strawberry”.
@@ II Methods
We propose empowering a text-based LLM to act as a versatile instruction follower for speech recognition in multi-talker speech scenarios. The proposed method consists of three major components: a large language model as the foundational model fine-tuned with low-rank adaptation (LoRA) [32], dual speech encoders with corresponding adapters, and training data construction. We denote the proposed model as MT-LLM in the subsequent sections.
@@ II-C Task Descriptions
To validate the performance of MT-LLM in executing speech recognition based on instructions in multi-talker scenarios, we simulated multi-talker overlapped speech audios from single-talker speech corpora, and designed six tasks along with corresponding instructions and generated the respective target text label. As illustrated in Fig. 1 (b), the text instructions pertain to the following tasks:
@@ III-A Training Data Construction
We simulated multi-talker speech audios from single-talker speech corpora to accommodate the versatile task instructions, following the protocol outlined in [18]. Specifically, the utterances are simulated primarily from the 960-hour LibriSpeech [35] training set, comprising mixtures of two or three talkers. The start time for each talker is randomly sampled, resulting in overlapped mixtures. Additionally, the 180-hour German subset of CoVoST 2333Note that CoVoST 2 is originally designed for speech translation, while we use its German subset for ASR in our study. [36], along with its corresponding German target text, is employed to mix with LibriSpeech utterances to support the Target-Lingual ASR task.
@@ III-C Model Settings and Training details
The proposed MT-LLM employs the encoder of Whisper-large-v2444https://huggingface.co/openai/whisper-large-v2 and WavLM-base [^5] https://huggingface.co/microsoft/wavlm-base to extract speech representations, with Llama-2-chat-7b as the backbone LLM. All parameters of above models are frozen. The LLM is fine-tuned using LoRA with a rank of 32. The adapters for both speech encoders consist of two convolution layers to down-sample and align the representations within the temporal domain, followed by a bottleneck adapter [39] and a linear layer. The outputs of both adapters have a time stride of 80 ms and a dimension of 2,048. The total number of parameters in MT-LLM is 7.55 billion, of which 1% (76.6 million) are trainable.
@@ IV-A Results on Multi-Talker ASR
We evaluate the single-talker ASR performance on the LibriSpeech and CoVoST 2 German test set, and multi-talker ASR performance on the 2- and 3-speaker LibriSpeechMix and the home-made En-De-Mixed test set. As shown in Table I, the proposed MT-LLM demonstrates promising results across all datasets, outperforming D2V-Sidecar-DB [28], a recent representative work employing PIT. SOT-Conformer [25] is a state-of-the-art model specifically designed and extensively trained for multi-talker ASR on the LibriSpeechMix dataset. On the 2-speaker LibriSpeechMix, MT-LLM achieves results comparable to SOT-Conformer, while lags behind on the 3-speaker set. Given that MT-LLM serves as an exploratory work aimed at exploring the potential of LLM for versatile instruction-based ASR in complex environments, rather than solely pushing the limits of multi-talker ASR, we argue that the gap is understandable. Despite being trained on both single-talker and multi-talker data, we note that SOT-Conformer falls short in single-talker scenarios, indicating its specialization in multi-talker ASR impairs its single-talker performance. In contrast, MT-LLM performs consistently good in single-talker ASR, demonstrating that its instruction-understanding capabilities help maintain high performance in simpler, single-talker settings. SALMONN [6] is a speech LLM trained on various speech-related tasks including overlapped speech recognition. However, using its official prompt, we observed significant inferior performance compared to our approach as shown in Table I.
We list wav2vec2-Large-XLSR-53-German model [^6] https://huggingface.co/jonatasgrosman/wav2vec2-large-xlsr-53-german performance on De for reference. Given that the audio samples containing German is relatively limited and noisier than LibriSpeech, MT-LLM still demonstrates satisfactory results on De and En-De-Mix test sets, showcasing the model’s ability to simultaneously distinguish and transcribe multiple languages from multi-talker speech.
@@ IV-B Results on Versatile Tasks Based on Instructions
We investigate the performance of MT-LLM on various ASR tasks described in Section II-C. The evaluations are conducted using the 2- and 3-speaker LibriSpeechMix and En-De-Mix test sets, with the results presented in Table II and Table III.
@@
The results in Table II and Table III highlight the effectiveness of the MT-LLM model across different instruction-based ASR tasks. In both the 2-speaker and 3-speaker mixed scenarios. MT-LLM achieves impressive performance for Target-Talker (TT), Keyword Tracing (KT), and Sex-Specific (SS) ASR tasks, which are within a reasonable range of the best-matching results. This suggests that MT-LLM is proficient at isolating and accurately transcribing speech from specified talkers based on instructions focusing on different speaker attributes. However, for Order-Specific (OS) ASR tasks, there is a gap compared to the best-matching result, indicating that MT-LLM slightly falls short in determining the order in which the talker appears. We anticipate significant improvements by using additional positional embedding for the speech part, employing larger speech modality adapters, or fully fine-tuning the speech encoder with heavier training as in [25]. Target-Lingual (TL) ASR task presents a more pronounced challenge for the model, due to that the mixed audios containing German are relatively limited and noisier compared to English set. We observe that performance of the TL task in both English and German is close, resulting in similar WER to best-matching situation. This indicates room for improvement through language-specific embeddings or higher-quality multi-lingual data.
@@ IV-C Ablation Study
First, we conduct an ablation study to examine the impact of using dual speech encoders. We train two models specifically on the multi-talker (MT) ASR task: one using dual speech encoders and the other using only the Whisper encoder. As shown in Table IV, incorporating the WavLM encoder into the architecture leads to notable performance improvements, underscoring the importance of multi-layer acoustic information captured by WavLM in improving speech recognition accuracy, particularly in three-talker scenarios.
We also investigate the effect of multi-task training on improving the basic MT ASR task. As illustrated in Table IV, training on a variety of tasks not only endows the system with the versatility to handle diverse ASR tasks but also improves performance on the basic MT ASR task. This indicates that the different tasks are interdependent and complementary, helping to supervise the model and enhance its overall speech comprehension capabilities.
@@ IV-D Limitations and Future Work
Despite MT-LLM’s promising performance on various ASR tasks in multi-talker scenarios, we acknowledge several limitations. First, the primary intention of this study is to explore the ability of LLMs to capture specific talkers according to instructions, and transcribe their speech in multi-talker scenarios. Therefore, MT-LLM is not designed to be a universal speech LLM for a broader spectrum of tasks. Second, the experiments are conducted on limited simulated datasets for insightful observation, rather than on real-world datasets, due to resource constraints. In the future, we anticipate the development of a more comprehensive speech LLM for cocktail party environments with meticulously crafted training data and schemes.
@@ V Conclusion
In this work, we present a pioneering exploration into the use of large language models (LLMs) for instruction-based speech recognition in multi-talker scenarios. We utilize the Whisper encoder to extract semantic context information and WavLM to capture multi-layer acoustic information indicating speaker characteristics, thereby enabling the foundational LLM to effectively handle speech modality input. With parameter-efficient fine-tuning, the proposed MT-LLM demonstrates remarkable capabilities in comprehending and transcribing speech based on versatile instructions related to multi-talker ASR, target talker ASR, and ASR based on specific talker attributes such as sex, occurrence order, language, and keyword spoken. Comprehensive experiments reveal promising performance in complex multi-talker environments, highlighting the potential of LLMs to enhance speech-related tasks and improve human-computer interaction in challenging settings.
VI Acknowledgments
This research is partially supported by the HKSARG Research Grants Council’s Theme-based Research Grant Scheme (Project No. T45-407/19N) and by the CUHK Stanley Ho Big Data Decision Analytics Research Centre.